# LLM Watermarking Using Mixtures and Statistical-to-Computational Gaps

### v1

@@ 4. Watermarking Scheme - Open Source Setting
Let’s describe the construction of the perturbation vector $\Delta\in\mathbb{R}^{d}$ in detail. First, pick a $k$-sparse subset $S\subset[d]$ uniformly at random, and compute the vector
@@
supported on $S$. Let $r$ be a Rademacher random variable and $G\sim N(0,I)$. For a fixed $\varepsilon>0$, define the perturbation vector
(4.1) $\Delta\coloneqq\begin{cases}\varepsilon(G+\mu),\quad\text{if $r=1$}\\ \varepsilon(G-\mu),\quad\text{if $r=-1$.}\end{cases}$
@@
(4.2) $q(i)\coloneqq\frac{e^{\ell(i)+\Delta(i)}}{e^{\ell(1)+\Delta(1)}+\cdots+e^{\ell(d)+\Delta(d)}},$
@@
  • $H_{0}:\Delta_{1},\ldots,\Delta_{N}$ was sampled from $N(0,\varepsilon^{2}I)$.
  • $H_{1}:$ $\Delta_{1},\ldots,\Delta_{N}$ was sampled from the mixture (4.1).
@@
In the sparsity regime $k\ll\sqrt{d}$, no polynomial-time algorithm on $N$ can solve (4) with less than $N=\tilde{\Omega}(k^{2})$ samples, assuming the $k$-BPC conjecture is true. On the other hand, there exists a computationally infeasible algorithm that solves (4) with $k=\Theta(k\log d)$ samples.
@@
The best the adversary can hope for is to predict $\Delta_{adv}\sim N(\mu_{adv},\varepsilon^{2}I)$, for some $\mu_{adv}$ satisfying $\|\mu-\mu_{adv}\|_{2}\geq 1/2$. (Indeed, if $\|\mu_{adv}-\mu\|_{2}<1/2$, then the adversary would solve the sparse mean estimation hypothesis testing which is not possible in polynomial time thanks to Theorem 4.1. Indeed, to learn the mean $\mu$ with higher accuracy one would need more than $\tilde{\Theta}(k^{2})\gg N$ samples.)
Notice that the total variation distance between the $\Delta_{adv}$ and $\Delta$ (or equivalently $-\mu$) is at least $\Phi(-1/4)>0.4$, where $\Phi(\cdot)$ is the cumulative density function of a standard multivariate Gaussian. ∎
@@
4.1. Watermark Detection - Open Source Setting
@@
The text $x_{1},\ldots,x_{N}$. The key: $G_{1},\ldots,G_{N}$ and $\varepsilon$. Tolerance $\delta$.
True or False. $Z\leftarrow\frac{1}{N}\sum_{j=1}^{N}G_{j}(x_{j})$.
if $Z\geq\varepsilon\sqrt{2\log(1/\delta)/N}$ then return: True
end ifreturn: False
Let $c_{0}$ be the absolute constant in [20, Proposition 2.7.1 from (b) to (e)]. Our main result for the open source setting is the following theorem:
Theorem 4.3 (Main Result for the Open Source Setting).
Let $c_{0}$ as above. The watermarking scheme described the rule (4.2) with $\varepsilon\leq 1/2$ and the detection Algorithm 2 satisfies the following properties:
  1. (1)
    It is unremovable.
Moreover, for any $\delta\in(0,1)$,
  1. (2)
    The watermarking scheme is sound with probability at least $1-\delta$.
  2. (3)
    Let $p^{*}_{j}$ denote the probability of the most likely token in the vocabulary at step $j=1,\ldots,N$. If, for some $\gamma>0$,
    (4.3) $\frac{\varepsilon^{2}}{1200}\frac{1}{N}\sum_{j=1}^{N}(1-p_{j}^{\ast})\geq\varepsilon\sqrt{\frac{2\log(1/\delta)}{N}}+\gamma.$
    Then, with probability at least $1-e^{-\gamma^{2}N/672c_{0}^{2}\varepsilon^{2}}$, the watermark scheme is complete.
@@
Before we proceed to the prove, we require some preliminary results. The proofs of the preliminary results are postponed to the Appendix. Recall that the softmax functions (2.1) and (4.2).
@@
${\mathbb{E}}G(x)\geq\frac{\varepsilon^{2}}{88e\sqrt{8\pi}}(1-p^{\ast}(1)).$
@@
We leave the proofs to the Appendix. To state our main result about the open source setting, let $c_{B}$ be the absolute constant in the sub-exponential Bernstein’s inequality [20, Theorem 2.8.1].
Proof.
We already proved the unremovability requirement in Proposition 4.2.
Soundness: Notice that under the null hypothesis, $x_{1},\ldots,x_{N}$ are independent from $G_{1},\ldots,G_{N}$, therefore the test $Z$ defined in Algorithm 2 is the distributed as $Z\sim N(0,\varepsilon^{2}/N)$. By the standard estimate for the Gaussian tail,
$\mathbb{P}\{Z\geq\varepsilon t\}\leq e^{-t^{2}N/2},$
which is at most $\delta$ for $t=\sqrt{2\log(1/\delta)/N}$.
Completness: Assume that the text is watermarked. Set $Z_{0}\coloneqq 0$ and notice that
$Z_{k}\coloneqq Z_{k-1}+G_{k}(x_{k})-{\mathbb{E}}\{G_{k}(x_{k})|x_{1},\ldots,x_{k-1}\},$
@@
$Y_{k}\coloneqq[Z_{k}-Z_{k-1}]|x_{1},\ldots,x_{k-1},$
are sub-exponential. In addition to this, [20, Proposition 2.7.1] implies that there exists a constant $c_{0}>0$ for which all the sub-exponential random variables $Y_{k}$ satisfy
${\mathbb{E}}e^{\lambda|Y_{k}|}\leq e^{\lambda^{2}c_{0}^{2}\|Y_{k}\|_{\psi_{1}}^{2}}\quad\text{for all}\quad|\lambda|\leq\frac{1}{c_{0}\|Y_{k}\|_{\psi_{1}}}.$
It follows from the sub-exponential version of the Azuma-Hoeffding’s inequality [21, Theorem 2.3] and Proposition 4.4 that
$\begin{split}&\mathbb{P}\bigg{\{}Z\geq\frac{\varepsilon^{2}}{88e\sqrt{8\pi}}\frac{1}{N}\sum_{j=1}^{N}(1-p_{j}^{\ast})-t\bigg{\}}\geq 1-e^{-t^{2}N/2c_{0}^{2}84\varepsilon^{2}}.\end{split}$
@@
5. Appendix
5.1. Proof of Proposition 4.4
Proof.
@@ 5. Appendix
${\mathbb{E}}\bigg{[}g_{k}\frac{e^{\Delta(k)+\ell(k)}}{e^{\ell(1)+\Delta(1)}+\cdots+e^{\ell(d)+\Delta(d)}}\bigg{]}\geq e^{-1/\sqrt{k}}{\mathbb{E}}\bigg{[}g_{k}\frac{e^{g_{k}}}{e^{\ell(1)+g_{1}-\ell(k)}+\cdots+e^{\ell(d)+g_{d}-\ell(k)}}\bigg{]}.$
Clearly, $e^{-1/\sqrt{k}}\geq e^{-1}$ as $k\geq 1$. Next, denote by ${\mathbb{E}}_{j}$ the expectation with respect to the randomness of $g_{j}$ only. By the iterated law of expectation, we compute the expectation term in the right-hand side by
@@
(5.1) $\frac{\varepsilon^{2}}{1.05\sqrt{8\pi}}\sum_{k=1}^{d}{\mathbb{E}}_{1,\ldots,k-1,k+1,\ldots,d}\bigg{[}\frac{\alpha_{k}}{(\alpha_{k}+1)^{2}}\bigg{]}.$
@@
(5.2) $\sum_{i\neq k}e^{\ell(i)-\ell(k)+g_{i}}\leq 4.55\sum_{i\neq k}e^{\ell(i)-\ell(k)}\quad\text{and}\quad\sum_{i\neq k}e^{\ell(i)-\ell(k)+g_{i}}\geq\frac{1}{4}\sum_{i\neq k}e^{\ell(i)-\ell(k)}.$
By the second estimate in (5.2), we have that
@@
Finally, notice that the first estimate in (5.2) implies that
@@
Since (5.1) is non-negative, we have that
@@
We conclude that $\mathbb{P}(\mathcal{E})\geq\min\{1/2-e^{-2},3/4-1/2\}=1/4$, which finishes the proof. ∎
5.2. Proof of Lemma 4.5
Proof.
@@
${\mathbb{E}}e^{|G(x)|/\tau}\leq e^{1/\sqrt{k}}\sum_{k=1}^{d}{\mathbb{E}}\bigg{[}e^{g_{k}/\tau}\frac{e^{g_{k}}}{e^{\ell(1)-\ell(k)+g_{1}}+\ldots+e^{\ell(d)-\ell(k)+g_{d}}}\bigg{]}\leq 2,$
@@
(5.3) $\begin{split}&{\mathbb{E}}\bigg{[}\frac{e^{g_{k}}}{\alpha_{k}+e^{g_{k}}}\bigg{]}=\frac{1}{2}{\mathbb{E}}\bigg{[}\frac{e^{-|g_{k}|}}{\alpha_{k}+e^{-|g_{k}|}}\bigg{]}+\frac{1}{2}\bigg{[}\frac{e^{|g_{k}|}}{\alpha_{k}+e^{|g_{k}|}}\bigg{]}\\ &\geq\frac{1}{\alpha_{k}+1}\left(\frac{1}{2}{\mathbb{E}}e^{-|g_{k}|}+\frac{1}{2}\right)\\ &\geq\frac{1}{\alpha_{k}+1}\left(\frac{0.95}{2}e^{-2\varepsilon}+\frac{1}{2}\right)\quad\text{(as $\mathbb{P}\{|g_{k}|\geq 2\varepsilon\}\geq 0.95$ )}\\ &\geq\frac{0.67}{\alpha_{k}+1}\quad\text{(as $\varepsilon\leq 1/2$)}.\end{split}$
@@
(5.4) $\begin{split}&{\mathbb{E}}\bigg{[}\frac{e^{2g_{k}}}{(\alpha_{k}+e^{g_{k}})^{2}}\bigg{]}=\frac{1}{2}\left({\mathbb{E}}\bigg{[}\frac{e^{2|g_{k}|}}{(\alpha_{k}+e^{|g_{k}|})^{2}}\bigg{]}+{\mathbb{E}}\bigg{[}\frac{e^{-2|g_{k}|}}{(\alpha_{k}+e^{-|g_{k}|})^{2}}\bigg{]}\right)\\ &\leq\frac{1}{2}\left(2e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{e^{-2|g_{k}|}}{(\alpha_{k}+e^{-|g_{k}|})^{2}}\bigg{]}\right)\\ &=\frac{1}{2}\left(2e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{1}{(\alpha_{k}e^{|g_{k}|}+1)^{2}}\bigg{]}\right)\\ &=\left(e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{1}{2(\alpha_{k}e^{|g_{k}|}+1)^{2}}\bigg{]}\right)\\ &=\frac{1}{(\alpha_{k}+1)^{2}}\left(e^{2\varepsilon^{2}}+\frac{1}{2}\right)\\ &\leq\frac{1}{(\alpha_{k}+1)^{2}}(\sqrt{e}+\frac{1}{2})\quad\text{(as $\varepsilon\leq 1/2$)}.\end{split}$
Putting together (5.4) and (5.3) and the fact that $\varepsilon<1/2$, we obtain that

### v2

@@ 4.1 The watermarking scheme
Let’s describe the construction of the perturbation vector $\Delta\in\mathbb{R}^{d}$ in detail. First, pick a $k$-sparse subset $S\subset[d]$ uniformly at random, and compute the unit norm vector
@@
supported on $S$. Let $r$ be a Rademacher random variable. For a fixed $\varepsilon>0$, let $G\sim N(0,\varepsilon^{2}I)$ and define the perturbation vector
$\Delta\coloneqq\begin{cases}G+\varepsilon\mu,\quad\text{if $r=1$}\\ G-\varepsilon\mu,\quad\text{if $r=-1$.}\end{cases}$ (4.1)
@@
$q(i)\coloneqq\frac{e^{\ell(i)+\Delta(i)}}{e^{\ell(1)+\Delta(1)}+\cdots+e^{\ell(d)+\Delta(d)}},$ (4.2)
@@ 4.2 Unremovability
$\quad H_{0}$ $:\Delta_{1},\ldots,\Delta_{N}\sim N(0,\varepsilon^{2}I)$ (4.3)
[…]
@@
In the sparsity regime $k\ll\sqrt{d}$, no polynomial-time algorithm on $N$ can solve (4.3) with less than $N=\tilde{\Omega}(k^{2})$ samples, assuming the $k$-BPC conjecture is true. On the other hand, there exists a computationally infeasible algorithm that solves (4.3) with $k=\Theta(k\log d)$ samples.
@@
The best the adversary can hope for is to predict $\Delta_{adv}\sim N(\mu_{adv},\varepsilon^{2}I)$, for some $\mu_{adv}$ satisfying $\|\mu-\mu_{adv}\|_{2}\geq\varepsilon/2$. (Indeed, if $\|\mu_{adv}-\mu\|_{2}<\varepsilon/2$, then the adversary would solve the sparse mean estimation hypothesis testing which is not possible in polynomial time thanks to Theorem 4.1. Indeed, to learn the mean $\mu$ with higher accuracy one would need more than $\tilde{\Theta}(k^{2})\gg N$ samples.)
Notice that the total variation distance between the $\Delta_{adv}$ and $\Delta$ is at least $\Phi(-1/4)>0.4$, where $\Phi(\cdot)$ is the cumulative density function of a standard multivariate Gaussian. ∎
@@
4.3 Soudness and Completness
@@
The text $x_{1},\ldots,x_{N}$. The key: $G_{1},\ldots,G_{N}$. Tolerance $\delta$.
True or False. Compute $Z=\frac{1}{N}\sum_{j=1}^{N}G_{j}(x_{j})$. If $Z\geq\varepsilon\sqrt{2\log(1/\delta)/N}$ return True, else return False.
Proposition 4.3 (Soundness).
The watermarking scheme described in Section 4.1 is sound with probability at least $1-\delta$.
Proof.
Notice that under the null hypothesis, $x_{1},\ldots,x_{N}$ are independent from $G_{1},\ldots,G_{N}$, therefore the test $Z$ defined in Algorithm 2 is the distributed as $Z\sim N(0,\varepsilon^{2}/N)$. By the standard estimate for the Gaussian tail,
$\mathbb{P}\{Z\geq\varepsilon t\}\leq e^{-t^{2}N/2},$
which is at most $\delta$ for $t=\sqrt{2\log(1/\delta)/N}$. ∎
Let $c_{0}$ be the absolute constant in [22, Proposition 2.7.1 from (b) to (e)]. Our main result for the open source setting is the following theorem:
Theorem 4.4 (Completeness).
Let $\delta\in(0,1)$, $N\geq 2\log(1/\delta)$. Let $c_{0}$ as above and let $p^{*}_{j}$ denote the probability of the most likely token in the dictionary at step $j=1,\ldots,N$. If, for some $\gamma>0$ and $\varepsilon\leq 1/2$,
$\frac{\varepsilon^{2}}{1200}\frac{1}{N}\sum_{j=1}^{N}(1-p_{j}^{\ast})\geq\varepsilon\sqrt{\frac{2\log(1/\delta)}{N}}+\gamma.$ (4.4)
Then, with probability at least $1-e^{-\gamma^{2}N/672c_{0}^{2}\varepsilon^{2}}$, the watermark scheme described in Section 4.1 is complete.
@@
Before we proceed to the prove, we require some preliminary results. Recall the softmax functions (2.1) and (4.2).
@@
${\mathbb{E}}G(x)\geq\frac{\varepsilon^{2}}{1200}(1-p^{\ast}(1)).$
@@
We leave the proof of Proposition 4.5 to Appendix B and the proof of Lemma 4.6 to Appendix C.
Proof of Theorem 4.4.
Assume that the text is watermarked. Set $Z_{0}\coloneqq 0$ and notice that
$Z_{j}\coloneqq Z_{j-1}+G_{k}(x_{j})-{\mathbb{E}}\big{[}G_{j}(x_{j})|x_{1},\ldots,x_{j-1}\big{]},$
@@
$Y_{j}\coloneqq[Z_{j}-Z_{j-1}]|x_{1},\ldots,x_{j-1},$
are sub-exponential. In addition to this, [22, Proposition 2.7.1] implies that there exists a constant $c_{0}>0$ for which all the sub-exponential random variables $Y_{j}$ satisfy
${\mathbb{E}}e^{\lambda|Y_{j}|}\leq e^{\lambda^{2}c_{0}^{2}\|Y_{j}\|_{\psi_{1}}^{2}}\quad\text{for all}\quad|\lambda|\leq\frac{1}{c_{0}\|Y_{j}\|_{\psi_{1}}}.$
It follows from the sub-exponential version of the Azuma-Hoeffding’s inequality [23, Theorem 2.3] and Proposition 4.5 that
$\begin{split}&\mathbb{P}\bigg{\{}Z\geq\frac{\varepsilon^{2}}{1200}\frac{1}{N}\sum_{j=1}^{N}(1-p_{j}^{\ast})-t\bigg{\}}\geq 1-e^{-t^{2}N/2c_{0}^{2}84\varepsilon^{2}}.\end{split}$
@@
Appendix A Proof of Lemma 3.7
We start with the proof of the lower bound. By the Hoeffding-Chernoff inequality for the hyper-geometric distribution (see [9, Section 27.5]), for each $i=2,\ldots,d/10$, we have:
$\big{|}I\cap[10i]\big{|}\geq i\quad\big{|}I^{c}\cap[10i]\big{|}\geq i,$ (A.1)
with probability at least $1-2e^{-8i/5}$. Notice that for any $i$ for which the condition (A.1) holds, the monotonicity of $p$ implies that at least $i$ elements of $\{p(i)\}_{i\in I}$ are greater or equal to $p_{10i}$. By definition of $w$, this implies that $w_{i}\geq p_{10i}$. Similarly, we have $\overline{w}_{i}\geq p_{10i}$ which implies $\min\{w_{i},\overline{w}_{i}\}\geq p_{10i}$.
Since for each $i=2,\ldots,d/10$, the event (A.1) holds with probability at least $21-2e^{-8i/5}\geq 1/$ 2, we have
${\mathbb{E}}\min\{w_{i},\overline{w}_{i}\}\geq\frac{p_{10i}}{2}.$
Additionally, with probability $1/2$ we have $1\in I$ and $2\in I^{c}$ (vice-versa), so
${\mathbb{E}}\min\big{\{}w_{1},\overline{w}_{1}\big{\}}\geq\frac{p_{2}}{2}.$
We conclude that
${\mathbb{E}}\sum_{i=1}^{d/2}\min\{w(i),\overline{w}(i)\}\geq\frac{1}{2}\big{(}p_{2}+p_{10}+p_{20}+\cdots+p_{d}\big{)}$ (A.2)
To complete the full sum (from $i=2$ to $d$), notice that
$8p_{2}\geq p_{3}+\cdots+p_{9}\quad\text{and}\quad 10p_{10i}\geq p_{10i}+\cdots+p_{10i+9}.$ (A.3)
Combining (A.2) and (A.3), we obtain
${\mathbb{E}}\sum_{i=1}^{d/2}\min\{w_{i},\overline{w}_{i}\}\geq\frac{1}{20}\sum_{i=2}^{d}p_{i}=\frac{1}{20}(1-p_{1}).$
The proof of the upper bound is shorter. In fact, almost surely we have
$\begin{split}&\sum_{i=1}^{d}\min\big{\{}w_{i},\overline{w}_{i}\big{\}}\leq\min\bigg{\{}\sum_{i=1}^{d/2}w_{i},\sum_{i=1}^{d/2}\overline{w}_{i}\bigg{\}}\\ &\leq\min\bigg{\{}\sum_{i\in I}p_{i},\sum_{i\in I^{c}}p_{i}\bigg{\}}\leq\sum_{i=2}^{d}p_{i}.\end{split}$
where the last inequality holds because either $I$ or $I^{c}$ does not contain the element $1$. The proof is complete.
Appendix B Proof of Proposition 4.5
@@
${\mathbb{E}}\bigg{[}g_{k}\frac{e^{\Delta(k)+\ell(k)}}{e^{\ell(1)+\Delta(1)}+\cdots+e^{\ell(d)+\Delta(d)}}\bigg{]}\geq e^{-\varepsilon/\sqrt{k}}{\mathbb{E}}\bigg{[}g_{k}\frac{e^{g_{k}}}{e^{\ell(1)+g_{1}-\ell(k)}+\cdots+e^{\ell(d)+g_{d}-\ell(k)}}\bigg{]}.$
Clearly, $e^{-\varepsilon/\sqrt{k}}\geq e^{-1}$ as $k\geq 1$. Next, denote by ${\mathbb{E}}_{j}$ the expectation with respect to the randomness of $g_{j}$ only. By the iterated law of expectation, we compute the expectation term in the right-hand side by
@@
$\frac{\varepsilon^{2}}{1.05\sqrt{8\pi}}\sum_{k=1}^{d}{\mathbb{E}}_{1,\ldots,k-1,k+1,\ldots,d}\bigg{[}\frac{\alpha_{k}}{(\alpha_{k}+1)^{2}}\bigg{]}.$ (B.1)
@@
$\sum_{i\neq k}e^{\ell(i)-\ell(k)+g_{i}}\leq 4.55\sum_{i\neq k}e^{\ell(i)-\ell(k)}\quad\text{and}\quad\sum_{i\neq k}e^{\ell(i)-\ell(k)+g_{i}}\geq\frac{1}{4}\sum_{i\neq k}e^{\ell(i)-\ell(k)}.$ (B.2)
By the second estimate in (B.2), we have that
@@
Finally, notice that the first estimate in (B.2) implies that
@@
Since (B.1) is non-negative, we have that
@@
We conclude that $\mathbb{P}(\mathcal{E})\geq\min\{1/2-e^{-2},3/4-1/2\}=1/4$, which finishes the proof.
Appendix C Proof of Lemma 4.6
@@
${\mathbb{E}}e^{|G(x)|/\tau}\leq e^{\varepsilon/\sqrt{k}}\sum_{k=1}^{d}{\mathbb{E}}\bigg{[}e^{g_{k}/\tau}\frac{e^{g_{k}}}{e^{\ell(1)-\ell(k)+g_{1}}+\ldots+e^{\ell(d)-\ell(k)+g_{d}}}\bigg{]}\leq 2,$
@@
$\begin{split}&{\mathbb{E}}\bigg{[}\frac{e^{g_{k}}}{\alpha_{k}+e^{g_{k}}}\bigg{]}=\frac{1}{2}{\mathbb{E}}\bigg{[}\frac{e^{-|g_{k}|}}{\alpha_{k}+e^{-|g_{k}|}}\bigg{]}+\frac{1}{2}\bigg{[}\frac{e^{|g_{k}|}}{\alpha_{k}+e^{|g_{k}|}}\bigg{]}\\ &\geq\frac{1}{\alpha_{k}+1}\left(\frac{1}{2}{\mathbb{E}}e^{-|g_{k}|}+\frac{1}{2}\right)\\ &\geq\frac{1}{\alpha_{k}+1}\left(\frac{0.95}{2}e^{-2\varepsilon}+\frac{1}{2}\right)\quad\text{(as $\mathbb{P}\{|g_{k}|\geq 2\varepsilon\}\geq 0.95$ )}\\ &\geq\frac{0.67}{\alpha_{k}+1}\quad\text{(as $\varepsilon\leq 1/2$)}.\end{split}$ (C.1)
@@
$\begin{split}&{\mathbb{E}}\bigg{[}\frac{e^{2g_{k}}}{(\alpha_{k}+e^{g_{k}})^{2}}\bigg{]}=\frac{1}{2}\left({\mathbb{E}}\bigg{[}\frac{e^{2|g_{k}|}}{(\alpha_{k}+e^{|g_{k}|})^{2}}\bigg{]}+{\mathbb{E}}\bigg{[}\frac{e^{-2|g_{k}|}}{(\alpha_{k}+e^{-|g_{k}|})^{2}}\bigg{]}\right)\\ &\leq\frac{1}{2}\left(2e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{e^{-2|g_{k}|}}{(\alpha_{k}+e^{-|g_{k}|})^{2}}\bigg{]}\right)\\ &=\frac{1}{2}\left(2e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{1}{(\alpha_{k}e^{|g_{k}|}+1)^{2}}\bigg{]}\right)\\ &=\left(e^{2\varepsilon^{2}}\frac{1}{(\alpha_{k}+1)^{2}}+{\mathbb{E}}\bigg{[}\frac{1}{2(\alpha_{k}e^{|g_{k}|}+1)^{2}}\bigg{]}\right)\\ &=\frac{1}{(\alpha_{k}+1)^{2}}\left(e^{2\varepsilon^{2}}+\frac{1}{2}\right)\\ &\leq\frac{1}{(\alpha_{k}+1)^{2}}(\sqrt{e}+\frac{1}{2})\quad\text{(as $\varepsilon\leq 1/2$)}.\end{split}$ (C.2)
Putting together (C.2) and (C.1) and the fact that $\varepsilon<1/2$, we obtain that